3. La Besurta

’FORAU’ DE ’AIGUALLUT’ 40 minutes

The departure point is the end of the asphalt trail at La Besurta camping area. Follow a well-marked track along the bottom of the valley. After 20 minutes you will

get to the Pllau d’Aiguallut, a large prairie that ends abruptly at the Forau d’Aiguallut, which you will reach in half an hour.

At Aiguallut the water that comes down from the Aneto glaciars and the Barrancs valley, pours down in an abyss (the Forau) and reappears at the Val d’Aran, birth place of the Garona river, which disembogues near the city of Burdeos, in the Atlantic Ocean.